OpenStudy (anonymous):

tan^2xsin^2x=tanx-sin^2x

OpenStudy (debbieg):

I'm pretty sure that's supposed to be tan^2x-sin^2x on the RHS Try using the Fundamental Identity on the LHS, to replace sin^2x.... then multiply it out and see what falls into place. :)

OpenStudy (debbieg):

Fundamental ID: \(\Large \sin^2x+\cos^2x=1\) rearrange it to solve it for \(\sin^2x\), then use that on the LHS.

OpenStudy (anonymous):

Oh okay, this helps me a lot. Thanks (:

OpenStudy (debbieg):

You're welcome. :)